OpenStudy (butterflyhope):

Check: 42 - (-17) + 24 = 35. Thx :)

OpenStudy (shamim):

-(-1)=?

OpenStudy (butterflyhope):

?

OpenStudy (shamim):

1?

jimthompson5910 (jim_thompson5910):

42 - (-17) + 24 is not equal to 35

OpenStudy (butterflyhope):

No....I don't think so at least. My other answer would have been 83, because; 42 + -17 = 59. 59 + 24 = 83. I got 35 from: -17 + 24 = 7 42 - 7 + 35.

OpenStudy (shamim):

i m telling u abt a formula

OpenStudy (shamim):

so u know -(-17)=?

jimthompson5910 (jim_thompson5910):

42 - (-17) + 24 42 + 17 + 24 59 + 24 83 So, 42 - (-17) + 24 = 83

OpenStudy (butterflyhope):

Oh sorry, brain is melted by thinking today. xD

jimthompson5910 (jim_thompson5910):

shamim is saying that -(-17) = +17 ie, subtracting a negative is the same as adding a positive

OpenStudy (antonio_xx2):

u need to know ur math lol 42 - (-17) + 24 = 83